Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability question_Answer evaluation

Expert replies
by pharmxanthan » Fri Jun 11, 2010 1:03 pm
Q) A bag contains 3 red balls, 5 yellow balls, and 4 green balls. Two balls are drawn randomly. What is the probability that the balls drawn contain balls of different colors?

A) Number of ways to select 2 balls out of 12 balls = 12C2 = 66.
Number of ways to select two balls of different colors (red & yellow OR red & green OR yellow & green)= (3C1*5C1)+(3C1*4C1)+(5C1*4C1)= 47
Probability = 47/66 (Ans)

Is the approach correct?

Thanks!
Join the discussion
Source: — Problem Solving |

by Patrick_GMATFix » Fri Jun 11, 2010 1:06 pm
yep the approach is correct. Haven't checked the numbers

-Patrick
  • Ask me about tutoring.
Join the discussion

by Stuart@KaplanGMAT » Fri Jun 11, 2010 6:57 pm
pharmxanthan wrote:Q) A bag contains 3 red balls, 5 yellow balls, and 4 green balls. Two balls are drawn randomly. What is the probability that the balls drawn contain balls of different colors?

A) Number of ways to select 2 balls out of 12 balls = 12C2 = 66.
Number of ways to select two balls of different colors (red & yellow OR red & green OR yellow & green)= (3C1*5C1)+(3C1*4C1)+(5C1*4C1)= 47
Probability = 47/66 (Ans)

Is the approach correct?

Thanks!
A very similar way to do it would be to look at the 3 cases and use the probability formula.

First case: choose red first, non-red second.

Prob = 3/12 * 9/11 = 27/132

Second case: choose yellow first, non-yellow second.

Prob = 5/12 * 7/11 = 35/132

Third case: choose green first, non-green second.

Prob = 4/12 * 8/11 = 32/132

Prob of case 1 OR case 2 OR case 3:

27/132 + 35/132 + 32/132 = 94/132 = 47/66

Of course, we could also solve using the "1 minus" formula:

Probability (what you want) = 1 - Probability (what you don't want)

In this question, we don't want two of the same colour.

Prob red/red = 3/12 * 2/11 = 6/132
Prob yellow/yellow = 5/12 * 4/11 = 20/132
Prob green/green = 4/12 * 3/11 = 12/132

Prob double = (6+20+12)/132 = 38/132 = 19/66

Prob not double = 66/66 - 19/66 = 47/66

Or, using combinations as you did:

Number of ways to select 2 balls: 12C2 = 12*11/2 = 66

Number of ways to select 2 of the same colour: 3C2 + 4C2 + 5C2 = 3 + 6 + 10 = 19

Prob of selecting 2 of the same colour = 19/66
Prob of NOT selecting 2 of the same colour = 66/66 - 19/66 = 47/66

* * *

When practicing, it's good to think of all the different ways you can attack a question; that way, when it comes to Test Day, you'll know which approach works best for you.
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion